# Two Characterizations of Geometrically Infinite Actions on Gromov Hyperbolic Spaces

### v1

@@ 1. Introduction
Let $X$ be a proper Gromov hyperbolic space, on which a countable group $G$ acts properly by isometries. The action naturally extends to a convergence group action on the Gromov boundary $\partial X$, which decomposes $\partial X$ into the limit set $\Lambda G$ and the domain of discontinuity. Let $\Pi\subset\Lambda G$ be the set of parabolic points. The action of $G$ on $X$ is called geometrically finite if there exists a sufficiently separated $G$-invariant family of horoballs $\{B_{p}:p\in\Pi\}$ so that the action on the complement $\mathrm{CH}(\Lambda G)\setminus\bigcup_{p\in\Pi}B_{p}$ is cocompact, where $CH(\Lambda G)$ denotes the weak convex hull of $\Lambda G$ in $X$. In this case, $\Lambda G\setminus\Pi$ consists precisely of conical points. We refer to Section 2 for precise definitions.
@@
The corresponding result for Kleinian groups was proved by Bonahon [BON86] and was recently generalized to pinched negatively curved manifolds by Kapovich and Liu [KL19]. In contrast to these works, our approach relies only on elementary arguments in hyperbolic geometry and does not invoke the Margulis lemma.
@@
Suppose a group $G$ admits a non-elementary proper action on a proper Gromov hyperbolic space $X$. Then the action is geometrically infinite if and only if the set of non-conical points is uncountable.
We remark that the negatively pinched manifold case was due to Kapovich and Liu [KL20] after their work [KL19] on Theorem 1.1. To put into perspective, for finitely generated geometrically infinite Kleinian groups, the non‑conical limit set has maximal Hausdorff dimension whose precise value was computed by Bishop–Jones [BJ97], subsequent contributions [KL20], and recently completed in [MY25] in full general case. The question of whether Theorem 1.2 holds in this level of generality arose naturally in the course of that study.
Finally, let us conclude the Introduction with an open question. By a theorem of Yaman [YAM04] (and of Bowditch [BOW98] in the cocompact case), geometrically finite actions on Gromov hyperbolic spaces can be recovered from the dynamics on the boundary. More precisely, if a convergence group action is geometrically finite (i.e., its limit set consists only of conical points and bounded parabolic points) then the group admits a geometrically finite action on some proper Gromov hyperbolic space. Moreover, the induced action on the boundary coincides with the original convergence group action. The following question asks whether the assumption of bounded parabolic points could be weakend as follows.
Question 1.3.
Suppose a discrete group $G$ admits a convergence group action on a compact metrizable space with countably many non-conical limit points. Does $G$ admit a geometrically finite action on some Gromov hyperbolic space?
@@
    This is related to a question of Kapovich [KAP05, Problem 7] asking whether every convergence group action can be realized as the boundary action of a proper action on some Gromov hyperbolic space. If so, then a positive answer would follow from Theorem 1.2 together with Yaman’s theorem. Moreover, in [SUN19], Bin Sun proved that the action constructed in [BOW98] on a hyperbolic space contains WPD (i.e. weak proper discontinuity) elements.
@@
    The assumption that the action is a convergence group action is necessary. It is easy to give examples of non-relatively hyperbolic groups $G$ which admit a minimal action on a compact metriziable space $M$ so that all but countably many points $\xi\in M$ are conical points in the usual sense.
@@
    Indeed, any non-elementary action on a locally infinite tree $T$ gives raise to the desired action on the compact space $M=\partial T\cup T^{0}$, the union of the ends of the tree and the vertex set. We equip $M$ with Bowditch topology (since $T$ is a fine graph in sense of [BOW12]). It is easy to check that the ends of the tree are conical points in the above sense.
@@ 2. Preliminaries
In this section we introduce the preliminaries we shall use on convergence group actions (cf. [TUK94, BOW99]) and hyperbolic spaces (cf. [BH13]). In particular, we follow the account of [BOW12] on geometrically finite actions.
@@ 2.1. Convergence actions
Given such an action, the limit set $\Lambda H$ of a subgroup $H\subset G$ is defined as the set of all accumulation points of the $H$-orbits in $M$. The action is said to be non-elementary if $\Lambda G>2$. In this case, the limit set $\Lambda G$ is the unique minimal $G$-invariant closed subset; moreover, $\Lambda G$ is a perfect set (i.e., it has no isolated points) and therefore contains uncountably many points.
@@
Next, define $L=\bigcap\limits_{n=1}^{\infty}f^{-n}(M\setminus U)$. Observe that $f^{-1}(M\setminus U)\subset M\setminus\overline{U}\subset M\setminus U$. Similarly, we have $f^{-1}L=L$, hence $fL=L$.
@@
A point $\xi\in M$ is called conical if there exist a sequence $\{g_{n}\}$ in $G$ and distinct points $a,b\in M$ such that $g_{n}x\to b$ locally uniformly for all $x\in M\setminus\{\xi\}$, while $g_{n}\xi\to a$. A point $\xi\in M$ is parabolic if its stabilizer $G_{\xi}$ is infinite and its limit set satisfies $\Lambda G_{\xi}=\{\xi\}$. A parabolic point $\xi$ is bounded parabolic if the action of $G_{\xi}$ on $M\setminus\{\xi\}$ is cocompact.
Note that both conical and parabolic points lie in the limit set $\Lambda G$. We denote by $\Lambda^{c}G$ and $\Lambda^{nc}G$ the sets of conical and non-conical points in $\Lambda G$, respectively. In particular, the fixed points of any hyperbolic element are conical.
A convergence group action of $G$ on $M$ is geometrically finite if every point of $\Lambda G$ is either conical or bounded parabolic. Note that the induced action of $G$ on $\Lambda G$ is a convergence group action. In the literature, geometrically finite actions are usually required to be minimal, i.e., $\Lambda G=M$. We deliberately adopt this convention to remain consistent with the notion of a geometrically finite action on a hyperbolic space; see Lemma 2.10.
2.2. Geometrically finite actions
Let $\delta\geq 0$. A geodesic space $X$ is called $\delta$-hyperbolic if every geodesic triangle in $X$ is $\delta$-thin; that is, each side is contained in the $\delta$-neighborhood of the union of the other two sides. In the sequel, we always assume that $X$ is a proper geodesic $\delta$-hyperbolic space.
The set of asymptotic classes of geodesic rays in $X$ is called the Gromov boundary, denoted $\partial X$. We set $\overline{X}:=X\cup\partial X$, which forms a compactification of $X$ under the cone topology. Note that any two points $x,y\in\overline{X}$ can be joined by a geodesic. We denote by $[x,y]$ an (often arbitrary) choice of geodesic between $x$ and $y$. Given a geodesic $\gamma=[x,y]$ and $x^{\prime},y^{\prime}\in\gamma$, we denote the subpath of $\gamma$ between $x^{\prime},y^{\prime}$ by $x^{\prime},y^{\prime}]_{\gamma}$.
For three distinct points $x,y,z\in\overline{X}$, an ideal triangle $\triangle(xyz)$ is the union $x,y]\cup[y,z]\cup[z,x]$. In a $\delta$-hyperbolic space, every ideal triangle is $2\delta$-thin. Furthermore, in a $\delta$-hyperbolic space $X$, any quadrilateral is $2\delta$-thin, and any ideal quadrilateral is $4\delta$-thin. These facts are left as exercises for the interested reader.
@@ 2.2. Geometrically finite actions
for all $x\neq y\in\overline{X}$. This metric is called a visual metric based at $z$ with parameter $\varepsilon$. The topology induced on $\partial X$ by any visual metric coincides with the cone topology (see [BH13]).
Given a path $p$ in $X$, we denote the endpoints of $p$ by $p_{-},p_{+}$. Given $\lambda\geq 1,C\geq 0$, a path $p$ in $X$ is called a $(\lambda,C)$-quasi-geodesic if $\operatorname{Len}(q)\leq\lambda d(q_{-},q_{+})+C$ for every connected subpath $q$ of $p$. Let $L>0,\lambda\geq 1$ and $C\geq 0$. A path $p$ is called an $L$-local $(\lambda,C)$-quasi-geodesic if every subpath of $p$ of length $L$ is a $(\lambda,C)$-quasi-geodesic.
@@
For any $\lambda\geq 1,C\geq 0$, there exists $L_{0}>0,\lambda^{\prime}\geq 1$ and $C^{\prime}\geq 0$ with the following property. Given $L\geq L_{0}$, let $\gamma$ be an $L$-local $(\lambda,C)$-quasi-geodesic. Then $\gamma$ is a $(\lambda^{\prime},C^{\prime})$-quasi-geodesic.
@@
where $\mathcal{P}(A)$ denotes the set of all subsets of $A$. Note that $\pi_{A}(x)$ is non-empty, thanks to the proper assumption on $X$.
@@
Let $\gamma$ be a geodesic line in $X$ and $x,y\in X$ such that $d(x^{\prime},y^{\prime})>8\delta$ for some $x^{\prime}\in\pi_{\gamma}(x)$ and $y^{\prime}\in\pi_{\gamma}(y)$. Then there exists a subpath $\gamma^{\prime}$ of $[x,y]$ such that $d_{\text{Haus}}(\gamma^{\prime},[x^{\prime},y^{\prime}])\leq 8\delta$, where $d_{\text{Haus}}$ denotes the Hausdorff distance.
@@
and similarly $d(q,y^{\prime})\leq 6\delta$. Since the (ideal) quadrilateral $(p,x^{\prime},y^{\prime},q)$ is $2\delta$-thin, we have $d_{\text{Haus}}([p,q],[x^{\prime},y^{\prime}])\leq 8\delta$. ∎
The projection to a geodesic line $\gamma$ can be extended to $\pi_{\gamma}:\overline{X}\setminus\{\gamma_{-},\gamma_{+}\}\to\gamma$ as follows. Suppose $\{x_{n}\}$ is a sequence in $X$ which converges to $\xi\in\partial X$ and $y_{n}\in\pi_{\gamma}(x_{n})$. Then any limit point of $\{y_{n}\}$ lies in $\pi_{\gamma}(\xi)$. Note that $d([x_{n},x_{m}],\gamma)>8\delta$ for sufficently large $m,n>0$. By Lemma 2.4 we obtain that $d(y_{n},y_{m})\leq 8\delta$. Therefore, the extension map $\pi_{\gamma}:\overline{X}\setminus\{\gamma_{-},\gamma_{+}\}\to\gamma$ is $8\delta$-coarsely continuous. Hence Lemma 2.4 can be generalized as follows.
@@
Let $\gamma$ be a geodesic line in $X$ and $x,y\in\overline{X}$ such that $d(x^{\prime},y^{\prime})>24\delta$ for some $x^{\prime}\in\pi_{\gamma}(x)$ and $y^{\prime}\in\pi_{\gamma}(y)$. Then there exists a subpath $\gamma^{\prime}$ of $[x,y]$ such that $d_{\text{Haus}}(\gamma^{\prime},[x^{\prime},y^{\prime}])\leq 24\delta$, where $d_{\text{Haus}}$ denotes the Hausdorff distance.
@@
$d_{\text{Haus}}(\gamma_{1},\gamma_{2})\leq D$
for some $D>0$. Then $d_{\text{Haus}}(\pi_{\gamma_{1}}(x),\pi_{\gamma_{2}}(x))\leq 2D+100\delta$ for any $x\in\overline{X}$.
@@
$d_{\text{Haus}}([y^{\prime},x^{\prime}],[z,y_{2}]_{\gamma_{2}})\leq 24\delta.$
@@
We will use the following Morse Lemma for $\delta$-hyperbolic spaces.
Lemma 2.7.
If $\gamma$ is a $(\lambda,C)$-quasi-geodesic in $X$, then $d_{\text{Haus}}(\gamma,[\gamma_{-},\gamma_{+}])\leq B$.

### v2

@@ Abstract.
Key words and phrases:
Gromov hyperbolic space, geometrical finiteness, convergence group action
@@ 1. Introduction
Let $X$ be a proper Gromov hyperbolic space, on which a countable group $G$ acts properly by isometries. The action naturally extends to a convergence group action on the Gromov boundary $\partial X$, which decomposes $\partial X$ into the limit set $\Lambda G$ and the domain of discontinuity. Let $\Pi\subset\Lambda G$ be the set of parabolic points. The action of $G$ on $X$ is called geometrically finite if there exists a sufficiently separated $G$-invariant family of horoballs $\{B_{p}:p\in\Pi\}$ so that the action on the complement $\mathrm{CH}(\Lambda G)\setminus\bigcup_{p\in\Pi}B_{p}$ is cocompact, where $\mathrm{CH}(\Lambda G)$ denotes the weak convex hull of $\Lambda G$ in $X$. In this case, $\Lambda G\setminus\Pi$ consists precisely of conical points. We refer to Section 2 for precise definitions.
@@
The corresponding result for Kleinian groups was proved by Bonahon [BON86] and was recently generalized to pinched negatively curved manifolds by Kapovich and Liu [KL19]. In contrast to these works, our approach relies only on elementary arguments in hyperbolic geometry and does not invoke the Margulis lemma. In detail, we define a region that acts as a horoball (see Section 3.2) and use Lemma 2.3 to directly obtain a hyperbolic element of word length $2$, thus circumventing the essential obstacle in previous works.
@@
Suppose that a group $G$ admits a non-elementary proper action on a proper Gromov hyperbolic space $X$. Then the action is geometrically infinite if and only if the set of non-conical points is uncountable.
We remark that the Kleinain group case was done by Bishop [BIS96] and the pinched negatively curved manifold case was established by Kapovich and Liu [KL20] after their work [KL19] on Theorem 1.1. To put this into perspective, for finitely generated geometrically infinite Kleinian groups, the non‑conical limit set has maximal Hausdorff dimension; its precise value was computed by Bishop–Jones [BJ97], subsequent contributions by [KL20], and was recently completed in the full general case by [MY25]. The question of whether Theorem 1.2 holds at this level of generality arose naturally in the course of that study.
Remark 1.3.
As a direct corollary of Theorem 1.2, such an action is geometrically finite if and only if every point of $\Lambda G$ is either conical or parabolic. That is, we can remove the bounded assumption on parabolic points. See Definition 2.4.
Finally, let us conclude the Introduction with an open question. By a theorem of Yaman [YAM04] (and of Bowditch [BOW98] in the cocompact case), geometrically finite actions on Gromov hyperbolic spaces can be recovered from the dynamics on the boundary. More precisely, if a convergence group action is geometrically finite (i.e., its limit set consists only of conical points and bounded parabolic points) then the group admits a geometrically finite action on some proper Gromov hyperbolic space. Moreover, the induced action on the boundary coincides with the original convergence group action. In the same sprint of those works, one can ask whether the analogue of Theorem 1.2 holds in a purely dynamical term of convergence groups.
Question 1.4.
Suppose that a discrete group $G$ admits a convergence group action on a compact metrizable space with countably many non-conical limit points. Does $G$ admit a geometrically finite action on some Gromov hyperbolic space?
@@
    This is related to a question of Kapovich [KAP05, Problem 7] asking whether every convergence group action can be realized as the boundary action of a proper action on some Gromov hyperbolic space. If so, then a positive answer would follow from Theorem 1.2 together with Yaman’s theorem. It is also deserved mentioning that in [SUN19], Sun proved that the action constructed in [BOW98] on a hyperbolic space contains WPD (i.e., weak proper discontinuity) elements.
@@
    The assumption that the action is a convergence group action is necessary. It is easy to give examples of non-relatively hyperbolic groups $G$ which admit a minimal action on a compact metrizable space $M$ such that all but countably many points $\xi\in M$ are conical points in the usual sense.
@@
    Indeed, any non-elementary action on a locally infinite tree $T$ gives rise to the desired action on the compact space $M=\partial T\cup T^{0}$, the union of the ends of the tree and the vertex set. We equip $M$ with Bowditch topology (since $T$ is a fine graph in the sense of [BOW12]). It is easy to check that the ends of the tree are conical points in the above sense.
@@ 2. Preliminaries
In this section we introduce the preliminaries we will use on convergence group actions (cf. [TUK94, BOW99]) and hyperbolic spaces (cf. [BH13]). In particular, we follow the account of [BOW12] on geometrically finite actions.
@@ 2.1. Convergence actions
Definition 2.2.
Given such an action, the limit set $\Lambda H$ of a subgroup $H\subset G$ is defined as the set of all accumulation points of the $H$-orbits in $M$.
The action is said to be non-elementary if $\Lambda G>2$.
If the action is non-elementary, the limit set $\Lambda G$ is the unique minimal $G$-invariant closed subset; moreover, $\Lambda G$ is a perfect set (i.e., it has no isolated points) and therefore contains uncountably many points.
@@
Next, define $L=\bigcap\limits_{n=1}^{\infty}f^{-n}(M\setminus U)$. Observe that $f^{-1}(M\setminus U)\subset M\setminus\overline{U}\subset M\setminus U$. Similarly, we have $L\neq\varnothing$ and $f^{-1}L=L$, hence $fL=L$.
@@
Definition 2.4.
A point $\xi\in M$ is called conical if there exist a sequence $\{g_{n}\}$ in $G$ and distinct points $a,b\in M$ such that $g_{n}x\to b$ locally uniformly for all $x\in M\setminus\{\xi\}$, while $g_{n}\xi\to a$.
A point $\xi\in M$ is parabolic if its stabilizer $G_{\xi}$ is infinite and its limit set satisfies $\Lambda G_{\xi}=\{\xi\}$.
A parabolic point $\xi$ is bounded parabolic if the action of $G_{\xi}$ on $M\setminus\{\xi\}$ is cocompact.
Note that both conical and parabolic points lie in the limit set $\Lambda G$. We denote by $\Lambda^{c}G$ and $\Lambda^{nc}G$ the sets of conical and non-conical points in $\Lambda G$, respectively. In particular, the fixed points of any hyperbolic element are conical. The following definition has its origin in the work of Beardon-Maskit [BM74] for Kleinian groups.
Definition 2.5.
A convergence group action of $G$ on $M$ is geometrically finite if every point of $\Lambda G$ is either conical or bounded parabolic.
Note that the induced action of $G$ on $\Lambda G$ is a convergence group action. In the literature, geometrically finite actions are usually required to be minimal, i.e., $\Lambda G=M$. We deliberately adopt this definition to remain consistent with the notion of a geometrically finite action on a hyperbolic space; see Definition 2.22 and Lemma 2.23.
2.2. Gromov hyperbolic geometry
Definition 2.6.
Let $\delta\geq 0$. A geodesic space $X$ is called $\delta$-hyperbolic if every geodesic triangle in $X$ is $\delta$-thin; that is, each side is contained in the $\delta$-neighborhood of the union of the other two sides.
In the sequel, we always assume that $X$ is a proper geodesic $\delta$-hyperbolic space.
Definition 2.7.
The set of asymptotic classes of geodesic rays in $X$ is called the Gromov boundary, denoted by $\partial X$. We set $\overline{X}:=X\cup\partial X$, which forms a compactification of $X$ under the cone topology.
Note that any two points $x,y\in\overline{X}$ can be joined by a geodesic. We denote by $[x,y]$ a (often arbitrary) choice of geodesic between $x$ and $y$. Given a geodesic $\gamma=[x,y]$ and $x^{\prime},y^{\prime}\in\gamma$, we denote the subpath of $\gamma$ between $x^{\prime},y^{\prime}$ by $x^{\prime},y^{\prime}]_{\gamma}$.
For three distinct points $x,y,z\in\overline{X}$, an ideal triangle $\triangle(xyz)$ means the union $x,y]\cup[y,z]\cup[z,x]$ (with endpoints possibly in boundary). In a $\delta$-hyperbolic space, every ideal triangle is $2\delta$-thin. Furthermore, in a $\delta$-hyperbolic space $X$, any quadrilateral is $2\delta$-thin, and any ideal quadrilateral is $4\delta$-thin. These facts are left as exercises for the interested reader.
Definition 2.8.
@@ 2.2. Gromov hyperbolic geometry
for all $x\neq y\in\overline{X}$. This family of metrics are called visual metrics based at $z$ with parameter $\varepsilon$.
The topology induced on $\partial X$ by any visual metric coincides with the cone topology (see [BH13]).
Given a path $p$ in $X$, we denote the endpoints of $p$ by $p_{-},p_{+}$.
Definition 2.11.
Given $\lambda\geq 1,C\geq 0$, a path $p$ in $X$ is called a $(\lambda,C)$-quasi-geodesic if $\operatorname{Len}(q)\leq\lambda d(q_{-},q_{+})+C$ for every connected subpath $q$ of $p$.
Let $L>0,\lambda\geq 1$ and $C\geq 0$. A path $p$ is called an $L$-local $(\lambda,C)$-quasi-geodesic if every subpath of $p$ of length $L$ is a $(\lambda,C)$-quasi-geodesic.
We will use the following Morse Lemma for $\delta$-hyperbolic spaces.
Lemma 2.12.
If $\gamma$ is a $(\lambda,C)$-quasi-geodesic in $X$, then $d_{\mathrm{Haus}}(\gamma,[\gamma_{-},\gamma_{+}])\leq B$.
@@
For any $\lambda\geq 1,C\geq 0$, there exist $L_{0}>0$, $\lambda^{\prime}\geq 1$ and $C^{\prime}\geq 0$ with the following property. Given $L\geq L_{0}$, let $\gamma$ be an $L$-local $(\lambda,C)$-quasi-geodesic. Then $\gamma$ is a $(\lambda^{\prime},C^{\prime})$-quasi-geodesic.
The isometries of a Gromov hyperbolic space $X$ can be classified into three mutually exclusive types: elliptic, parabolic, and hyperbolic elements. The notion of hyperbolic elements plays a crucial role in the sequel.
Definition 2.14.
An isometry $g$ on $X$ is called hyperbolic if the translation length $\tau(g)$ of $g$ defined by
$\tau(g):=\inf_{x\in X}d(x,gx)$
is strictly positive.
Note that the translation length is invariant under conjugacy.
Definition 2.15.
A hyperbolic isometry $g$ has exactly two fixed points on the boundary defined as $g^{+}:=\lim\limits_{n\to\infty}g^{n}o\in\partial X$ and $g^{-}:=\lim\limits_{n\to\infty}g^{-n}o\in\partial X$.
By an axis of $g$ we mean a geodesic line $g^{+},g^{-}]$ between the two fixed points.
By Lemma 2.12, there exists a constant $D_{0}>0$ depending only on $\delta$ such that the Hausdorff distance between any two axes of $g$ is bounded by $D_{0}$. The following result is well-known. We include a proof as we are unable to find appropriate reference.
Lemma 2.16.
There exists a constant $D_{1}>0$ such that
$\tau(g)\leq d(x,gx)\leq\tau(g)+D_{1}$
for any hyperbolic element $g\in G$ and any point $x$ on some axis of $g$.
Proof.
Suppose $\gamma$ is an axis of $g$. By definition, we have $d(x,gx)\geq\tau(g)$ for any $x\in\gamma$, and there exists $y\in X$ such that $\tau(g)\leq d(y,gy)\leq\tau(g)+\delta$. Then for any $n\in\mathbb{Z}$ and any $y^{\prime}\in g^{n}[y,gy]$, we have $gy^{\prime}\in g^{n+1}[y,gy]$, and
$\tau(g)\leq d(y^{\prime},gy^{\prime})\leq d(y^{\prime},g^{n+1}y)+d(gy^{\prime},g^{n+1}y)=d(y,gy)\leq\tau(g)+\delta.$
If $d([y,gy],\gamma)\leq 8\delta$, then there exist $z\in\gamma$ and $y^{\prime}\in[y,gy]$ such that
$d(z,gz)\leq d(z,y^{\prime})+d(y^{\prime},gy^{\prime})+d(gy^{\prime},gz)\leq\tau(g)+17\delta.$
If $d([y,gy],\gamma)>8\delta$, then $\operatorname{diam}(\pi_{\gamma}(y)\cup\pi_{\gamma}(gy))\leq 8\delta$ by Lemma 2.18 from below. Let $z$ be a point in $\pi_{\gamma}(y)$. Then $gz\in\pi_{g\gamma}(gy)$. By Lemma 2.20 below, we have $d_{\mathrm{Haus}}(\pi_{g\gamma}(gy),\pi_{\gamma}(gy))\leq 2D_{0}+100\delta$. Hence
$d(z,gz)\leq\operatorname{diam}(\pi_{\gamma}(y)\cup\pi_{\gamma}(gy))+d_{\mathrm{Haus}}(\pi_{g\gamma}(gy),\pi_{\gamma}(gy))\leq 2D_{0}+108\delta.$
In summary, there exists a point $z\in\gamma$ such that $d(z,gz)\leq\tau(g)+2D_{0}+200\delta$.
Analogously, for any $n\in\mathbb{Z}$ and any $z^{\prime}\in g^{n}[z,gz]$, we have
$\tau(g)\leq d(z^{\prime},gz^{\prime})\leq\tau(g)+2D_{0}+108\delta.$
For $n\in\mathbb{Z}$, choose $x_{n}\in\pi_{\gamma}(g^{n}z)$. Then $d(x_{n},g^{n}z)\leq D_{0}$ and hence $\lim_{n\to+\infty}x_{n}=g^{+},\lim_{n\to-\infty}x_{n}=g^{-}$. This implies that $\gamma=\bigcup_{n\in\mathbb{Z}}[x_{n},x_{n+1}]_{\gamma}$.
Therefore, given $x\in\gamma$, there exists $n\in\mathbb{Z}$ such that $x\in[x_{n},x_{n+1}]_{\gamma}$. Since the quadrilateral $(x_{n},g^{n}z,g^{n+1}z,x_{n+1})$ is $2\delta$-thin, for any $x\in[x_{n},x_{n+1}]_{\gamma}$ we have
$d(x,[x_{n},g^{n}z]\cup[g^{n}z,g^{n+1}z]\cup[g^{n+1}z,x_{n+1}])\leq 2\delta.$
If $d(x,[x_{n},g^{n}z]\cup[x_{n+1},g^{n+1}z])\leq 2\delta$, then either $d(x,g^{n}z)\leq D_{0}+2\delta$ or $d(x,g^{n+1}z)\leq D_{0}+2\delta$. Without loss of generality, assume $d(x,g^{n}z)\leq D_{0}+2\delta$. Then we have
$d(x,gx)\leq d(x,g^{n}z)+d(g^{n}z,g^{n+1}z)+d(g^{n+1}z,gx)\leq\tau(g)+4D_{0}+112\delta.$
If $d(x,[g^{n}z,g^{n+1}z])\leq 2\delta$, then there exists $z^{\prime}\in[g^{n}z,g^{n+1}z]$ such that $d(x,z^{\prime})\leq 2\delta$. Hence
$d(x,gx)\leq d(x,z)+d(z,gz)+d(gz,gx)\leq\tau(g)+2D_{0}+112\delta.$
Thus, taking $D_{1}=4D_{0}+112\delta$ completes the proof. ∎
Until the end of this subsection, we discuss the strong contracting property of the shortest projection to quasigeodesics. This is also standard but less written in the literature, so we give some details at the convenience of the reader.
Definition 2.17.
@@
where $\mathcal{P}(A)$ denotes the set of all subsets of $A$.
Note that $\pi_{A}(x)$ is non-empty, thanks to the proper assumption on $X$.
@@
Let $\gamma$ be a geodesic line in $X$ and $x,y\in X$ such that $d(x^{\prime},y^{\prime})>8\delta$ for some $x^{\prime}\in\pi_{\gamma}(x)$ and $y^{\prime}\in\pi_{\gamma}(y)$. Then there exists a subpath $\gamma^{\prime}$ of $[x,y]$ such that $d_{\mathrm{Haus}}(\gamma^{\prime},[x^{\prime},y^{\prime}])\leq 8\delta$, where $d_{\mathrm{Haus}}$ denotes the Hausdorff distance.
@@
and similarly $d(q,y^{\prime})\leq 6\delta$. Since the (ideal) quadrilateral $(p,x^{\prime},y^{\prime},q)$ is $2\delta$-thin, we have $d_{\mathrm{Haus}}([p,q],[x^{\prime},y^{\prime}])\leq 8\delta$. ∎
The projection to a geodesic line $\gamma$ can be extended to the boundary as follows:
$\pi_{\gamma}:\overline{X}\setminus\{\gamma_{-},\gamma_{+}\}\to\gamma$
Indeed, suppose $\{x_{n}\}$ is a sequence in $X$ which converges to $\xi\in\partial X$ and $y_{n}\in\pi_{\gamma}(x_{n})$. Then any limit point of $\{y_{n}\}$ lies in $\pi_{\gamma}(\xi)$. Note that $d([x_{n},x_{m}],\gamma)>8\delta$ for sufficiently large $m,n>0$. By Lemma 2.18 we obtain that $d(y_{n},y_{m})\leq 8\delta$. Therefore, the extension map $\pi_{\gamma}:\overline{X}\setminus\{\gamma_{-},\gamma_{+}\}\to\gamma$ is $8\delta$-coarsely continuous. Hence Lemma 2.18 can be generalized as follows.
@@
Let $\gamma$ be a geodesic line in $X$ and $x,y\in\overline{X}$ such that $d(x^{\prime},y^{\prime})>24\delta$ for some $x^{\prime}\in\pi_{\gamma}(x)$ and $y^{\prime}\in\pi_{\gamma}(y)$. Then there exists a subpath $\gamma^{\prime}$ of $[x,y]$ such that $d_{\mathrm{Haus}}(\gamma^{\prime},[x^{\prime},y^{\prime}])\leq 24\delta$, where $d_{\mathrm{Haus}}$ denotes the Hausdorff distance.
@@
$d_{\mathrm{Haus}}(\gamma_{1},\gamma_{2})\leq D$
for some $D>0$. Then $d_{\mathrm{Haus}}(\pi_{\gamma_{1}}(x),\pi_{\gamma_{2}}(x))\leq 2D+100\delta$ for any $x\in\overline{X}$.
@@
$d_{\mathrm{Haus}}([y^{\prime},x^{\prime}],[z,y_{2}]_{\gamma_{2}})\leq 24\delta.$
@@
2.3. Geometrically finite actions